# Entropic balance with feedback control: information equalities and tight inequalities

### v1

@@ Abstract
We consider overdamped physical systems evolving under a feedback-controlled fluctuating potential and in contact with a thermal bath at temperature $T$. A Markovian description of the dynamics, which keeps only the last value of the control action, is advantageous—both from the theoretical and the practical side—for the entropy balance. Novel second-law equalities and bounds for the extractable work are obtained, the latter being both tighter and easier to evaluate than those in the literature based on the whole chain of controller actions. The Markovian framework also allows us to prove that the bound for the extractable work that incorporates the unavailable information is saturated in a wide class of physical systems, for error-free measurements. These results are illustrated in a model system. For imperfect measurements, there appears an interval of measurement uncertainty, including the point at which work ceases to be extracted, where the new Markovian bound is tighter than the unavailable information bound.
@@
Feedback systems—from theoretical constructs like Maxwell’s demon [18] or Szilard’s engine [19] to natural systems such as molecular motors—challenge the reconciliation of the second law of thermodynamics with the role of information. The coarse-graining involved in modelling the controller—or “demon”—makes it necessary to rewrite the second law to include an information term linking system and demon variables [20, 21, 22, 6, 23, 24, 25, 26, 27, 12, 28, 29, 30, 13, 31].
@@
For a given, fixed, protocol, the utility of these second-law variants depends on how easily the associated information function can be computed and how tight the bound on extractable work is. In some cyclic information engines, like Szilard’s, a simplifying common feature is that system and controller states become independent at the end of each cycle. This reduces the thermodynamic balance to the single-measurement case, where mutual information is key [23]. However, in more complex systems like feedback flashing ratchets, system and controller states remain correlated across cycles [26]. One approach is to model the controller as having an infinite memory that stores all the control history, where the transfer entropy is key [6, 24, 26, 12, 7, 10].
@@
Our theoretical framework is a Markovian description of the joint stochastic process of system and controller [35], where only the most recent control value is kept—even in the presence of correlations between measurements. The Markovian entropy balance offers both theoretical and practical advantages. It yields a new information bound for extractable work that is both tighter and easier to evaluate than the bound based on the full control history [6, 24, 26, 12, 7, 10]. Also, the Markovian framework allows us to prove that the second-law inequality involving the so-called unavailable information [13] becomes an equality for a broad class of systems. All these results are illustrated using a model system, including a comparison between the new Markovian bound and the unavailable information bound.
@@
General Markovian framework– We consider a general class of physical systems, described by overdamped Brownian motion with friction $\gamma$ in a thermal bath at temperature $T=\beta^{-1}$  [^3] We take $k_{B}=1$ throughout the paper.. The system state $x$ evolves under the action of a conservative force, deriving from a fluctuating potential $V(x,c)$, and an external force $f$; the total force is thus $F(x,c)=-\partial_{x}V(x,c)+f$. The variable $c$ describes the fluctuation of the potential, $V(x,c)\neq V(x,c^{\prime})$ if $c\neq c^{\prime}$, and is set by a feedback mechanism. In the following, we summarise the key points of the Markovian framework that are necessary for our analysis, see [35] and Sec. I of [37] for more details.
@@
Figure 1: Sketch of the system dynamics. The feedback controller measures the system state at regular times $t_{n}=n\Delta t_{\text{m}}$, $x_{n}\equiv x(t_{n})$, $n=0,1,\ldots$ The value of the controller variable $c_{n}$ in the time interval $\tau_{\text{FP}}^{(n+1)}\equiv(t_{n}^{+},t_{n+1}^{-})$ is drawn from a probability distribution $\Theta(c_{n}|x_{n})$, conditioned to the measured $x_{n}$. The system follows overdamped Fokker-Planck (FP) evolution under a fluctuating potential $V(x,c_{n})$, different for each value of $c_{n}$, in the same interval. Then, up to time $t_{k}^{+}$, a chain of control actions $\vec{c}_{k}=\{c_{0},c_{1},\ldots,c_{k}\}$ is generated.
[…]
@@
The feedback controller updates its state in the infinitesimal time interval $\tau_{\text{m}}^{(n)}\equiv(t_{n}^{-},t_{n}^{+})$: it takes the value $c_{n}$ with conditional probability $\Theta(c_{n}|x_{n})$ at $t_{n}^{+}$. Then, (i) $P(x,c,t_{n}^{+})=\Theta(c|x)P(x,t_{n})$, and (ii) $P(x,t)$ is continuous at $t_{n}$, $P(x,t_{n}^{+})=P(x,t_{n}^{-})$. For error-free measurement, $x_{n}$ determines univocally the next controller value $c_{n}$: $\mathcal{D}$ splits into non-overlapping subsets $\mathcal{X}_{c}$, $\cup_{c}\mathcal{X}_{c}=\mathcal{D}$, and the potential $V(x,c_{n})$ is on if $x_{n}\in\mathcal{X}_{c_{n}}$. Then, $\Theta(c|x)=\bm{1}_{c}(x)$, where $\bm{1}_{c}(x)$ is the indicator function of the interval $\mathcal{X}_{c}$. If there are errors in the measurement, $\Theta(c|x)$ takes values different from zero or unity—see the numerical section for a specific example.
@@
The $H$-theorem for the above class of systems implies that a time-periodic state with period $\Delta t_{\text{m}}$ is reached for long times, $n\geq n_{R}\gg 1$. We shift the time origin, $\{n\to n-n_{R},t\to t-n_{R}\Delta t_{\text{m}}\}$, so that $n\geq 0$ and $t\geq 0$ correspond to the long-time regime. From now on, $P(x,c,t)$ thus refers to the long-time periodic solution; $P(x,t)$ and $P(c)$ refer to the marginal distributions—the latter is time-independent. We denote the chains of measured positions and corresponding control actions up to time $t_{n}^{+}$ by $\vec{x}_{n}=\{x_{0},x_{1},\ldots,x_{n}\}$ and $\vec{c}_{n}=\{c_{0},c_{1},\ldots\,c_{n}\}$, respectively, with joint probability $P(\vec{x}_{n},\vec{c}_{n})$. Interestingly, $\vec{x}_{n}$ is a Markov process but $\vec{c}_{n}$ is not: $P(x_{n}|\vec{x}_{n-1})=P(x_{n}|x_{n-1})$ but $P(c_{n}|\vec{c}_{n-1})\neq P(c_{n}|c_{n-1})$.
@@
Thermodynamic balance– It is in the long-time periodic state that we formulate the thermodynamic balance. Let us consider the change of the average value of any quantity $A(x,c)$, $\left\langle A\right\rangle(t)\equiv\sum_{c}\int dx\,A(x,c)P(x,c,t)$, over the interval $(t_{0}^{+},t_{k}^{+})$ depicted in Fig. 1, with $k$ cycles. Periodicity entails: (i) the values of $\left\langle A\right\rangle$ at the initial and final instants of the interval $\tau_{\text{FP}}^{(n)}$, $t_{n-1}^{+}$ and $t_{n}^{-}$ are independent of $n$, thus we denote them by $\left\langle A\right\rangle^{+}$ and $\left\langle A\right\rangle^{-}$, respectively, and (ii) its variation $\Delta\!\left\langle A\right\rangle$ in any period $(t_{n-1}^{+},t_{n}^{+})=\tau_{\text{FP}}^{(n)}\cup\tau_{\text{m}}^{(n)}$ is also $n$-independent and vanishes, $\Delta\!\left\langle A\right\rangle=0$. Moreover, $\Delta\!\left\langle A\right\rangle$ can be split into its increments $\Delta_{\text{FP}}\left\langle A\right\rangle$ for the FP evolution in the interval $\tau_{\text{FP}}^{(n)}$ and $\Delta_{\text{m}}\left\langle A\right\rangle$ at the $n$-th control update in the infinitesimal interval $\tau_{\text{m}}^{(n)}$: $\Delta\!\left\langle A\right\rangle=\Delta_{\text{FP}}\left\langle A\right\rangle+\Delta_{\text{m}}\left\langle A\right\rangle=0$, $\Delta_{\text{FP}}\left\langle A\right\rangle\equiv\left\langle A\right\rangle^{-}-\left\langle A\right\rangle^{+}$, $\Delta_{\text{m}}\left\langle A\right\rangle\equiv\left\langle A\right\rangle^{+}-\left\langle A\right\rangle^{-}$.
@@
where $S_{c|x}(t)=\sum_{c}\int dx\,P(x,c,t)\ln P(c|x,t)$. Moreover, for $J_{\text{s}}(x|c)\neq 0$, $\Sigma_{xc}$ can be split into positive-definite housekeeping and excess contributions [44, 45, 46, 47, 48, 28],
@@
For open-loop control, $I_{+}=0$ and thus $\Delta_{\text{m}}\,I=-I_{-}<0$; for error-free closed-loop control, $I_{+}=S_{c}$ and $I_{-}=S_{c}-S_{c|x}^{-}$, i.e. $\Delta_{\text{m}}\,I=S_{c|x}^{-}>0$.
Equation (7) is the first key result of our Letter, which must be compared with the existing inequalities based on the full history of control actions [6, 24, 26, 12, 7, 10],
@@
which is derived within our framework in Sec. III of [37]. Therein, $\overline{\mathcal{I}_{\vec{c}}}\geq 0$ is the average transfer entropy of the chain of control actions over $k$ cycles, with $k\to\infty$,
$k\,\overline{\mathcal{I}_{\vec{c}}}=\!\!\sum_{n=1}^{k}\mathcal{I}_{\vec{c}}^{(n)},\;\;\mathcal{I}_{\vec{c}}^{(n)}=-\!\sum_{\vec{c}_{n}}P(\vec{c}_{n})\ln P(c_{n}|\vec{c}_{n-1}).$ (9)
Note that $P(c_{n}|\vec{c}_{n-1})\!\neq\!P(c_{n}|c_{n-1})$: $\vec{c}_{n}$ is not Markovian. In Sec. IV of [37], we prove that $\overline{\mathcal{I}_{\vec{c}}}\geq\Delta_{\text{m}}I$. Then, the bounds in Eq. (7) are always tighter than those in (8):
@@
Below, we give the main steps of the derivation of Eq. (11); the full proof can be found in Sec. V of [37]. We start from the definition of $\mathcal{I}_{\vec{x}}$, analogous to Eq. (9), with the substitution $\vec{c}\to\vec{x}$. At variance with $\vec{c}$, $\vec{x}$ is a Markov process, i.e. $P(x_{n}|\vec{x}_{n-1})=P(x_{n}|x_{n-1})$, and
$\mathcal{I}_{\vec{x}}^{(n)}$ $=-\int dx_{n-1}dx_{n}P(x_{n-1},x_{n})\ln P(x_{n}|x_{n-1}).$ (12)
For error-free measurements, $\Theta(c_{i}|x_{i})=\bm{1}_{c_{i}}(x_{i})$, and $P(x_{n}|x_{n-1})=K(x_{n},\Delta t_{\text{m}}|x_{n-1},c_{n-1})$ if $x_{n-1}\in\mathcal{X}_{c_{n-1}}$, where $K(x,t|x^{\prime},c)$ is the propagator of the FP equation for the potential $V(x,c)$. Consequently, we get
@@
Now we turn our attention to $\mathcal{I}_{u}$, for which we have to introduce the reverse process [13]. In each trajectory of the forward process, the system evolution generates—through measurement and its associated feedback mechanism—chains $\vec{x}_{k}$ and $\vec{c}_{k}$; $(x_{0}\to c_{0})\longrightarrow(x_{1}\to c_{1})\longrightarrow(x_{2}\to c_{2})\cdots\longrightarrow(x_{k}\to c_{k})$. In each trajectory of the reverse process, time is inverted but there is no feedback: in each interval $\tau_{\text{FP}}^{(n)}$, the system moves under the same potential of the forward process, determined by $c_{n-1}$. Then, we have a fixed reverse chain of control actions, $\vec{c}_{k}^{R}\equiv\{c_{0}^{R}=c_{k}\to c_{1}^{R}=c_{k-1}\to\cdots c_{k}^{R}=c_{0}\}$. Also, we define reverse chains of system states $\vec{x}_{n}^{R}\equiv\{x_{0}^{R}=x_{k}\to x_{1}^{R}=x_{k-1}\to\cdots x_{n}^{R}=x_{k-n}\}$, generated with $\vec{c}_{k}^{R}$, with $n\leq k$. We denote the transition probability of reaching position $x_{n}^{R}$ under the protocol fixed by the forward process, given a prior chain of positions $\vec{x}_{n-1}^{R}$, by $P^{R}(x_{n}^{R}|\vec{x}_{n-1}^{R})$. In our notation, we write $\mathcal{I}_{u}^{(n)}=-\left\langle\ln P^{R}(x_{n}^{R}|\vec{x}_{n-1}^{R})\right\rangle$, i.e.
@@
For $f\neq 0$, detailed balance is not verified in general and Eq. (11) ceases to be valid. A generalisation thereof can be done in the limit $\Delta t_{\text{m}}\to\infty$, where the propagator of the Fokker-Planck equation tends to the non-equilibrium stationary distribution. By incorporating the housekeeping contribution to the entropy production, one gets—see Sec. VI of [37] for the proof,
@@
Application to a model information engine– Inspired by a class of information engines in which the feedback controller tries to decrease the energy of the particle and output work is done by the controller at the measurements [11, 12, 13, 14, 15, 16, 17], we consider a model with a potential $U(x)$ that can be inverted by the controller, V(x,c)=cU(x)V(x,c)=c\,U(x) with $c=\pm 1$. We aim to extract work at the measurements at the cost of possibly doing work during the FP evolution for $f\neq 0$. The interesting regime for an information engine is $\left\langle W\right\rangle=\left\langle W\right\rangle_{\text{m}}+\left\langle W\right\rangle_{\text{FP}}<0$.
@@
For the error-free case, $\Theta(c|x)=H(-c\,U(x))$, with $H(\cdot)$ being the Heaviside step function. Specifically, we have considered the piecewise potential $U(x)=V_{0}(|x|-L/2)/L$ in the interval [L,L][-L,L] with periodic boundary conditions (left panel of Fig. 2). When the particle position is measured with precision $\Delta x$, $\Theta(c|x)$ is proportional to the length of the interval $(x-\Delta x/2,x+\Delta x/2)$ with the “correct” sign of $U(x)$ (right panel of Fig. 2).
@@
Figure 2: (Left) Sketch of an information engine. The potential may be inverted by the feedback controller after measuring the position of the particle $x_{n}$ at times $t_{n}=n\Delta t_{\text{m}}$. In the error-free case, the inversion is done if the energy of the particle is decreased and work is always extracted. (Right) Introduction of error in the measurement. The particle position is measured with precision $\Delta x$. The function $\Theta(1|x)$, which gives the conditional probability of having the potential $V(x,c=1)$ at $t_{n}^{+}$ is plotted for different values of $\Delta x$, from $\Delta x=0$ (error-free measurement, perfect closed-loop control) to $\Delta x=1$ (maximum error, open-loop control).
@@
All numerical results are in dimensionless variables; by taking $L$ as the length unit, $T$ as the energy unit, and $\gamma L^{2}/T$ as the time unit. In the left panel of Fig. 3, we show the average work $\left\langle W\right\rangle$ and its bounds in Eq. (10) as a function of $\Delta x$ for a non-equilibrium situation with $f=0.7$. As predicted by our theoretical framework, the Markovian bounds are always tighter than their chain counterparts. It is worth noting that no net work can be extracted from the system, $\left\langle W\right\rangle>0$, for large enough error, $\Delta x>\Delta x_{0}(f)$; for $\Delta x=1$, the control becomes open-loop. The value $\Delta x_{0}(f)$ is quite accurately predicted by the Markovian bounds, but not by the chain bounds. This is further illustrated in the right panel of Fig. 3, where a phase diagram of the engine in the $(\Delta x,f)$ plane is plotted.
@@
Figure 3: (Left) Average work and bounds as a function of the uncertainty in the measurement. The data shown correspond to non-equilibrium situation with $f\neq 0$, specifically $f=0.7$. The behaviour for $f=0$ is qualitatively similar, eliminating the curves with the vanishing housekeeping contribution $\Sigma_{xc}^{\text{hk}}$. (Right) Phase diagram of the information engine in the $(\Delta x,f)$ plane. The regions $\left\langle W\right\rangle<0$ and $\left\langle W\right\rangle>0$ are separated by a line $\Delta x_{0}(f)$ (solid), which is accurately predicted by the Markovian bounds (black) but not by the chain bounds (blue). Additional parameters are $\Delta t_{\text{m}}=0.5$ and $V_{0}=5$.

### v2

@@ Abstract
We consider overdamped physical systems evolving under a feedback-controlled fluctuating potential and in contact with a thermal bath at temperature $T$. A Markovian description of the dynamics, which keeps only the last value of the control action, is advantageous—both from the theoretical and the practical side—for the entropy balance. Novel second-law equalities and bounds for the extractable work are obtained, the latter being both tighter and easier to evaluate than those in the literature based on the whole chain of controller actions. The Markovian framework also allows us to prove that the bound for the extractable work that incorporates the unavailable information is saturated in a wide class of physical systems, for error-free measurements. These results are illustrated in model systems. For imperfect measurements, there appears an interval of measurement uncertainty, including the point at which work ceases to be extracted, where the new Markovian bound is tighter than the unavailable information bound.
@@
Feedback systems—from theoretical constructs like Maxwell’s demon [18] or Szilard’s engine [19] to natural systems such as molecular motors—challenge the reconciliation of the second law of thermodynamics with the role of information. The coarse-graining involved in modelling the controller—or “demon”—makes it necessary to rewrite the second law to include an information term linking system and demon variables [20, 21, 22, 6, 23, 24, 25, 26, 27, 12, 28, 29, 30, 31, 13, 32].
@@
For a given protocol, the utility of these second-law variants depends on how easily the associated information function can be computed and how tight the bound on extractable work is. In some cyclic engines, like Szilard’s, a simplifying feature is system and controller states becoming independent at the end of each cycle. This reduces the thermodynamic balance to a single measurement, with mutual information being key [23]. Yet, in more complex systems like feedback flashing ratchets, system and controller states remain correlated across cycles [26]. One approach is to model the controller as having an infinite memory that stores all the control history, where transfer entropy is key [6, 24, 26, 12, 29, 7, 10].
@@
Our theoretical framework is a Markovian description of the system and controller joint stochastic process [36], where only the most recent control value is kept—even for correlated measurements. The Markovian entropy balance offers both theoretical and practical advantages. It yields a new information bound for extractable work that is both tighter and easier to evaluate than the bound based on the full control history [6, 24, 26, 12, 29, 7, 10]. Also, the Markovian framework allows us to prove that the second-law inequality involving the so-called unavailable information [13] becomes an equality for a broad class of systems. All results are illustrated using model systems, including a comparison between the new Markovian bound and the unavailable information bound.
@@
General Markovian framework– We consider a general class of physical systems, described by overdamped Brownian motion with friction $\gamma$ in a thermal bath at temperature $T=\beta^{-1}$  [^3] We take $k_{B}=1$ throughout the paper.. The system state $x$ evolves under the action of a conservative force, deriving from a fluctuating potential $V(x,c)$, and an external force $f$; the total force is thus $F(x,c)=-\partial_{x}V(x,c)+f$. The variable $c$ describes the fluctuation of the potential, $V(x,c)\neq V(x,c^{\prime})$ if $c\neq c^{\prime}$, and is set by a feedback mechanism  [^4] The variable $c$ may be multidimensional, we only assume that its set of possible values is countable.. In the following, we summarise the key points of the Markovian framework that are necessary for our analysis, see [36] and Sec. I of [39] for more details.
@@
Figure 1: Sketch of the system dynamics. The feedback controller measures the system state at regular times $t_{n}=n\Delta t_{\text{m}}$, $x_{n}\equiv x(t_{n})$, $n=0,1,\ldots$ The value of the controller variable $c_{n}$ in the time interval $\tau_{\text{FP}}^{(n+1)}\equiv(t_{n}^{+},t_{n+1}^{-})$ is drawn from a probability distribution $\Theta(c_{n}|x_{n})$, conditioned to the measured $x_{n}$. The system follows overdamped Fokker-Planck (FP) evolution under a fluctuating potential $V(x,c_{n})$, different for each value of $c_{n}$, in the same interval.
Figure 1 sketches the system evolution. Between measurements: (i) in the trajectory picture, the control variable equals the value $c_{n}$ stemming from the last measurement $x_{n}$, (ii) in the ensemble picture, the joint probability $P(x,c,t)$ obeys the FP equation with potential $V(x,c)$ [40, 41]: $\partial_{t}P(x,c,t)=-\partial_{x}J(x,c,t)$, with $\gamma J(x,c,t)\equiv-F(x,c)\,P(x,c,t)+\beta^{-1}\,\partial_{x}P(x,c,t)$. Since the control variable is frozen between measurements, the conditional probability $P(x|c,t)$ also obeys the FP equation with current $J(x|c,t)$. Its normalised stationary distribution and current are $P_{\text{s}}(x|c)$ and $J_{\text{s}}(x|c)$. The system variable $x$ is $d$-dimensional, $x\in\mathcal{D}\subset\mathbb{R}^{d}$—typically, $\mathcal{D}$ is either $\mathbb{R}^{d}$, with natural boundary conditions at infinity, or a cube with periodic boundary conditions. For $d>1$, $x$, $\partial_{x}$, $F$, $J$ are vectors and $\partial_{x}J$ is the divergence of $J$.
@@
The controller updates its state in the infinitesimal interval $\tau_{\text{m}}^{(n)}\equiv(t_{n}^{-},t_{n}^{+})$, $c_{n}$ with conditional probability $\Theta(c_{n}|x_{n})$ at $t_{n}^{+}$. Then, $P(x,c,t_{n}^{+})=\Theta(c|x)P(x,t_{n})$, and $P(x,t)$ is continuous at $t_{n}$, $P(x,t_{n}^{+})=P(x,t_{n}^{-})$. For error-free measurement, $x_{n}$ univocally determines $c_{n}$: $\mathcal{D}$ splits into non-overlapping subsets $\mathcal{X}_{c}$, $\cup_{c}\mathcal{X}_{c}=\mathcal{D}$, and the controller switches on $V(x,c_{n})$ if $x_{n}\in\mathcal{X}_{c_{n}}$. Then, $\Theta(c|x)=\bm{1}_{c}(x)$, where $\bm{1}_{c}(x)$ is the indicator function of the interval $\mathcal{X}_{c}$. The numerical section shows a specific example of $\Theta(c|x)$ for measurements with errors.
@@
Due to the periodic update of the feedback controller, a time-periodic state with period $\Delta t_{\text{m}}$ is reached for long times, $n\geq n_{R}\gg 1$. Redefining $\{n\to n-n_{R},t\to t-n_{R}\Delta t_{\text{m}}\}$, $n\geq 0$ and $t\geq 0$ correspond to the long-time regime. Henceforth, $P(x,c,t)$ refers to the long-time periodic solution; $P(x,t)$ and $P(c)$ to the marginal distributions—the latter is time-independent. The chains of measured positions and corresponding control actions up to time $t_{n}^{+}$ are denoted by $\vec{x}_{n}=\{x_{0},x_{1},\ldots,x_{n}\}$ and $\vec{c}_{n}=\{c_{0},c_{1},\ldots\,c_{n}\}$, with joint probability $P(\vec{x}_{n},\vec{c}_{n})$. Interestingly, $\vec{x}_{n}$ is Markovian but $\vec{c}_{n}$ is not:
$P(x_{n}|\vec{x}_{n-1})\!=\!P(x_{n}|x_{n-1}),\,P(c_{n}|\vec{c}_{n-1})\!\neq\!P(c_{n}|c_{n-1}).$ (1)
@@
Thermodynamic balance– We formulate the thermodynamic balance in the long-time periodic state. Consider the change in the average value of any quantity $A(x,c)$, $\left\langle A\right\rangle(t)\equiv\sum_{c}\int dx\,A(x,c)P(x,c,t)$, over the interval $(t_{0}^{+},t_{k}^{+})$ depicted in Fig. 1, with $k$ cycles. Periodicity entails $n$-independence of: (i) the values of $\left\langle A\right\rangle$ at the beginning and the end of $\tau_{\text{FP}}^{(n)}$, at $t_{n-1}^{+}$ and $t_{n}^{-}$, so we denote them by $\left\langle A\right\rangle^{+}$ and $\left\langle A\right\rangle^{-}$, respectively, and (ii) its variation $\Delta\!\left\langle A\right\rangle$ in any period $\tau_{\text{FP}}^{(n)}\cup\tau_{\text{m}}^{(n)}$, which vanishes, $\Delta\!\left\langle A\right\rangle=0$. Moreover, $\Delta\!\left\langle A\right\rangle$ can be split into its increments $\Delta_{\text{FP}}\left\langle A\right\rangle$ for the FP evolution in the interval $\tau_{\text{FP}}^{(n)}$ and $\Delta_{\text{m}}\left\langle A\right\rangle$ at the $n$-th control update in the infinitesimal interval $\tau_{\text{m}}^{(n)}$: $\Delta\!\left\langle A\right\rangle=\Delta_{\text{FP}}\left\langle A\right\rangle+\Delta_{\text{m}}\left\langle A\right\rangle=0$, $\Delta_{\text{FP}}\left\langle A\right\rangle\equiv\left\langle A\right\rangle^{-}-\left\langle A\right\rangle^{+}$, $\Delta_{\text{m}}\left\langle A\right\rangle\equiv\left\langle A\right\rangle^{+}-\left\langle A\right\rangle^{-}$.
@@
where $S_{c|x}(t)=\sum_{c}\int dx\,P(x,c,t)\ln P(c|x,t)$.
The steady distribution $P_{\text{s}}(x|c)$ may correspond to: (i) an equilibrium state, $J_{\text{s}}(x|c)=0$; (ii) a non-equilibrium steady state (NESS), $J_{\text{s}}(x|c)\neq 0$. For periodic boundary conditions, (i) arises for $f=0$ and (ii) arises for $f\neq 0$  [^6] For spatially periodic potentials, the force $f$ breaks detailed balance. $J_{\text{s}}(x|c){\color[rgb]{0,0,0}\neq 0}$ is independent of $x$ for $d=1$, whereas it may depend on $x$ for $d>1$.. In case (ii), $\Sigma_{xc}$ splits into positive-definite housekeeping and excess contributions [45, 46, 47, 48, 49, 28],
@@
For equilibrium, $\Sigma_{xc}^{\text{hk}}=0$: all dissipation stems from feedback and its associated information flow. For a NESS, $\Sigma_{xc}^{\text{hk}}>0$: additional dissipation stems from the housekeeping entropy production for sustaining the NESS [12].
Equation (8) is our first key result. The relevant regime for information engines is $\left\langle W\right\rangle=\left\langle W\right\rangle_{\text{m}}+\left\langle W\right\rangle_{\text{FP}}<0$. Two limit cases are worth considering: (i) open-loop control; measurement is not used for control, thus $c_{n}$ is independent of $x_{n}$ and $S_{c}=S_{c|x}$; $\Delta_{\text{m}}\,I=-I_{-}<0$, (ii) error-free closed-loop control; $c_{n}$ is enslaved to $x_{n}$ and $S_{c|x}=0$; $\Delta_{\text{m}}\,I=S_{c|x}^{-}>0$. Thus, feedback-control—maybe imperfect—is essential for extracting work.
The novel inequalities (8) for the extractable work must be compared with existing inequalities based on the full history of control actions [6, 24, 26, 12, 7, 10, 29],
@@
which is derived within our framework in Sec. III of [39]. Therein, $\overline{\mathcal{I}_{\vec{c}}}$ is the average transfer entropy of the chain of control actions over $k$ cycles, with $k\to\infty$,
$k\,\overline{\mathcal{I}_{\vec{c}}}=\!\!\sum_{n=1}^{k}\mathcal{I}_{\vec{c}}^{(n)},\;\;\mathcal{I}_{\vec{c}}^{(n)}=-\left[S_{x|\vec{c}}(t_{n}^{+})-S_{x|\vec{c}}(t_{n}^{-})\right],$ (10)
where $S_{x|\vec{c}}$ is defined as in Eq. (3), with $c\to\vec{c}$. Equation (10) involves calculating the entropy of a very long non-Markovian chain $\vec{c}_{n}$—recall Eq. (1777For error-free feedback control, Eq. (10) reduces to $\overline{\mathcal{I}_{\vec{c}}}=S_{\vec{c}}(t_{n}^{+})-S_{\vec{c}}(t_{n}^{-})=S_{c_{n}|\vec{c}_{n-1}}\geq 0$, see also [10].. In Sec. IV of [39], we prove that $\overline{\mathcal{I}_{\vec{c}}}\geq\Delta_{\text{m}}I$. Then, the bounds in Eq. (8) are always tighter than those in (9):
@@
Below, we give the main steps of the derivation of Eq. (12); the full proof can be found in Sec. V of [39]. From the definition $\mathcal{I}_{\vec{x}}^{(n)}\equiv S_{\vec{x}_{n}}-S_{\vec{x}_{n-1}}$, we have
$\mathcal{I}_{\vec{x}}^{(n)}$ $=-\int dx_{n-1}dx_{n}P(x_{n-1},x_{n})\ln P(x_{n}|x_{n-1}),$ (13)
where we have used the Markovianity of $\vec{x}_{n}$—as given by Eq. (1). For error-free measurements, $\Theta(c_{i}|x_{i})=\bm{1}_{c_{i}}(x_{i})$ and $P(x_{n}|x_{n-1})=K(x_{n},\Delta t_{\text{m}}|x_{n-1},c_{n-1})$ if $x_{n-1}\in\mathcal{X}_{c_{n-1}}$, where $K(x,t|x^{\prime},c)$ is the propagator of the FP equation for the potential $V(x,c)$. Consequently,
@@
Now we look into $\mathcal{I}_{u}$, for the reverse process [13]. In each trajectory of the forward process, chains $\vec{x}_{k}$ and $\vec{c}_{k}$ are generated through measurement and feedback: $(x_{0}\to c_{0})\longrightarrow(x_{1}\to c_{1})\longrightarrow(x_{2}\to c_{2})\cdots\longrightarrow(x_{k}\to c_{k})$—recall Fig. 1. In each trajectory of the reverse process, time is inverted but there is no feedback: in each interval $\tau_{\text{FP}}^{(n)}$, the system moves under the same potential of the forward process, determined by $c_{n-1}$. Then, we have a fixed reverse chain of control actions, $\vec{c}_{k}^{R}\equiv\{c_{0}^{R}=c_{k}\to c_{1}^{R}=c_{k-1}\to\cdots c_{k}^{R}=c_{0}\}$. Also, we define reverse chains of system states $\vec{x}_{n}^{R}\equiv\{x_{0}^{R}=x_{k}\to x_{1}^{R}=x_{k-1}\to\cdots x_{n}^{R}=x_{k-n}\}$, generated with $\vec{c}_{k}^{R}$, $n\leq k$. We denote the transition probability of reaching position $x_{n}^{R}$ under the protocol fixed by the forward process, given a prior chain of positions $\vec{x}_{n-1}^{R}$, by $P^{R}(x_{n}^{R}|\vec{x}_{n-1}^{R})$. In our notation, we write $\mathcal{I}_{u}^{(n)}=-\left\langle\ln P^{R}(x_{n}^{R}|\vec{x}_{n-1}^{R})\right\rangle$, i.e.
@@
For $f\neq 0$, detailed balance is not verified in general. Then, Eq. (12) ceases to be valid but it can be generalised for $\Delta t_{\text{m}}\to\infty$, where the FP propagator tends to the NESS distribution. By incorporating the housekeeping entropy production $\Sigma_{xc}^{\text{hk}}$, one gets—see Sec. VI of [39]:
@@
Application to a model information engine– Our framework applies to a broad class of information engines, including any colloidal particle undergoing overdamped motion (negligible inertia) in a fluctuating potential controlled by a Markovian (negligible delay) feedback device, as in Refs. [11, 14, 15, 16, 17]. In this work, we focus on engines where the feedback controller aims at reducing the particle’s energy. The goal is to maximize work extraction at the measurements [11, 12, 13, 14, 15, 16, 17], even at the cost of performing work during the Fokker–Planck evolution when $f\neq 0$.
@@
Inspired by the experimental setup in Ref. [11], we consider a model with a potential $U(x)$ that can be inverted by the controller, V(x,c)=cU(x)V(x,c)=c\,U(x) with $c=\pm 1$. Specifically, we have considered the piecewise potential $U(x)=V_{0}(|x|-L/2)/L$ in the interval [L,L][-L,L] with periodic boundary conditions (left panel of Fig. 2888This choice preserves the physical image of a sinusoidal potential [11] while simplifying analytical calculations.. For the error-free case, $\Theta(c|x)=H(-c\,U(x))$, with $H(\cdot)$ being the Heaviside step function. When the particle position is measured with precision $\Delta x$, $\Theta(c|x)$ is proportional to the length of the interval $(x-\Delta x/2,x+\Delta x/2)$ with the “correct” sign of $U(x)$ (right panel of Fig. 2).
@@
Figure 2: (Left) Sketch of an information engine. The potential may be inverted by the feedback controller after measuring the position of the particle $x_{n}$ at times $t_{n}=n\Delta t_{\text{m}}$. In the error-free case, the inversion is done—as shown by the arrows—if the potential energy decreases, so work is always extracted. For the sake of clarity, more than one period is shown. (Right) Introduction of error in the measurement. The particle position is measured with precision $\Delta x$. The function $\Theta(1|x)$, which gives the conditional probability of having the potential $V(x,c=1){\color[rgb]{0,0,0}=+U(x)}$ at $t_{n}^{+}$ is plotted for different values of $\Delta x$, from $\Delta x=0$ (error-free, perfect closed-loop control) to $\Delta x=1$ (maximum error, open-loop control).
@@
All numerical results are in dimensionless variables: $L$ is the length unit, $T$ is the energy unit, and $\gamma L^{2}/T$ is the time unit. In the left panel of Fig. 3, we show the average work $\left\langle W\right\rangle$ and its bounds in Eq. (11) as a function of $\Delta x$ for a non-equilibrium situation with $f=0.7$. As predicted by our theory, the Markovian bounds are always tighter than their chain counterparts. It is worth noting that no net work can be extracted from the system, $\left\langle W\right\rangle>0$, for large enough error, $\Delta x>\Delta x_{0}(f)$; for $\Delta x=1$, the control becomes open-loop. The value $\Delta x_{0}(f)$ is quite accurately predicted by the Markovian bounds, but not by the chain bounds. This is further illustrated in the right panel of Fig. 3, where a phase diagram of the engine in the $(\Delta x,f)$ plane is plotted.
@@
Figure 3: (Left) Average work and bounds as a function of the uncertainty in the measurement. The data shown correspond to non-equilibrium situation with $f\neq 0$, specifically $f=0.7$. The behaviour for $f=0$ is qualitatively similar, eliminating the curves with the vanishing housekeeping contribution $\Sigma_{xc}^{\text{hk}}$. (Right) Phase diagram of the information engine in the $(\Delta x,f)$ plane. The regions $\left\langle W\right\rangle<0$ and $\left\langle W\right\rangle>0$ are separated by a line $\Delta x_{0}(f)$ (solid), accurately predicted by the Markovian bounds (dashed black) but not by the chain ones (blue). Additional parameters are $\Delta t_{\text{m}}=0.5$ and $V_{0}=5$.